### v1

@@ Abstract
A nonlinear generalisation of the PageRank problem involving the Moore-Penrose inverse of the incidence matrix is developed for local graph partitioning purposes. The Levenberg-Marquardt method with a full rank Jacobian variant provides a strategy for obtaining a numerical solution to the generalised problem. Sets of vertices are formed according to the ranking supplied by the solution, and a conductance criterion decides upon the set that best represents the cluster around a starting vertex. Experiments on both synthetic and real-world inspired graphs demonstrate the capability of the approach to not only produce low conductance sets, but to also recover local clusters with an accuracy that consistently surpasses state-of-the-art algorithms.
@@ 1 Introduction
Graphs are a particularly convenient apparatus for representing associations and their value between objects. It is the very abstraction inherent in the definition of a graph that makes them so widely applicable. Vertices can be any kind of object (images [10], geographical locations [28], etc.), and this has no bearing on the graph at all. Furthermore, no restrictions are put on the number of edges, that is, associations, between any one vertex and the others.
@@
A graph may well exhibit a granular structure. Vertices that share a common property can be uncovered by a process called clustering. The resulting distinct groupings of vertices are referred to as communities or clusters. For example, if all associations are equally valuable, then it may be that communities are distinguished by the relatively high number of intra-community edges compared to the inter-community number of edges.
@@
Spectral clustering is a prominent algorithm that casts clustering as a graph partitioning problem. Clusters are determined through an analysis of the combinatorial Laplacian matrix eigenvectors [24, 36]. A generalisation of spectral clustering utilising the $p$-Laplacian [1, 35] was pursued in [5]. Improved clustering assignments were demonstrated in the bi-partitioning case. Recursive extension was, however, used for obtaining a greater number of clusters. Direct multiway approaches include approximating the $p$-orthogonality constraint [22, 27], the effective $p$-resistance [33], and exploiting concepts from total variation used commonly in image processing [4].
@@
When only a single cluster around a vertex or vertices is of interest, then an alternative strategy can be employed. Local clustering, as the name suggests, seeks to leverage local structure and information to identify a cluster. This has the potential to improve algorithmic run-time by avoiding computations over the entirety of a graph.
@@
Treatment of clustering as a partitioning problem carries over into local clustering. A cut can be found using a variation of the PageRank problem [2]. Since a PageRank vector provides a ranking of vertices (refer to [19] for information on functional rankings), sets of vertices can then be assembled according to their order and interrogated to reveal a cut. Just as with spectral clustering, the $p$-norm has found its way into local clustering algorithms. There is the nonlinear $p$-norm cut algorithm [21], and an algorithm based on the idea of diffusion with $p$-norm network flow [11].
@@
This work takes the system of linear equations that forms the PageRank problem as the starting point in the construction of a local clustering algorithm. Only clusters around a single vertex are considered. A modification to the PageRank problem is studied in Section 3. This is followed with the theoretical motivation for a nonlinear generalisation (that is, of the modified PageRank problem) involving the $p$-norm in Section 4. The Moore-Penrose inverse of the incidence matrix plays a very important role in the generalisation. It is shown in Section 4 that the generalisation can reduce to a system of linear equations closely resembling the modified PageRank problem. However, this is dependent on a number of conditions, with a major one being that the number of vertices approaches infinity. Additionally, an insight into the effect of the generalised problem on the cluster criterion based on an infinitesimal perturbation argument can be found in Section 4. The Levenberg-Marquardt method with a full rank Jacobian variant is presented in Section 5 for obtaining a numerical solution to the generalised problem.
@@
The next section is a short compendium of theories required for the development of the generalised problem and the identification of a cluster. Section 6 highlights the capability of the proposed algorithm on a number of synthetic and real-world inspired graphs. The algorithm performs strongly on a consistent basis, achieving results better than state-of-the-art algorithms.
@@ 2.1 Notions of graph theory
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set $\mathcal{V}$ of vertices and a set of unordered non-repeated pairs of distinct elements of $\mathcal{V}$ called edges, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Any two vertices $u,v\in\mathcal{V}$ are said to be adjacent or neighbours if they form an edge $e=\{u,v\}$ of $\mathcal{E}$. Their relationship is indicated by $u\sim v$. By definition, there can be no loops, i.e., $\{v,v\}\notin\mathcal{E}$, and only one edge can ever join two vertices. Such a graph is often described as being simple.
If a path exists between every pair of vertices $u$ and $v$ in $\mathcal{G}$, that is, a sequence of vertices $u=u_{1},u_{2},\ldots,u_{n},u_{n+1}=v$ with $e_{i}=\{u_{i},u_{i+1}\}\in\mathcal{E}$ for $i=1,2,\ldots,n$, then the graph is referred to as connected.
A graph $\mathcal{F}=(\dot{\mathcal{V}},\dot{\mathcal{E}})$ with $\dot{\mathcal{V}}\subseteq\mathcal{V}$ and $\dot{\mathcal{E}}\subseteq\mathcal{E}$ is said to be a subgraph of $\mathcal{G}$. Meanwhile, $\mathcal{G}$ becomes a host graph of $\mathcal{F}$. If $\mathcal{F}\neq\mathcal{G}$, then the subgraph is called proper. For $\mathcal{F}$ to be an induced subgraph, $\dot{\mathcal{E}}$ must contain all the edges of $\mathcal{E}$ required to maintain the adjacency relationships in host graph $\mathcal{G}$ between the vertices of $\dot{\mathcal{V}}$.
@@ 2.2 Conductance and sweep-cut
For a subset $\mathcal{S}$ of the vertices in a graph $\mathcal{G}$, i.e., $\mathcal{S}\subseteq\mathcal{V}$, the volume of $\mathcal{S}$ has the form
@@
The complement of $\mathcal{S}$ is $\bar{\mathcal{S}}=\mathcal{V}\setminus\mathcal{S}$. An edge boundary $\partial(\mathcal{S})$ of $\mathcal{S}$ collects the edges with one vertex in $\mathcal{S}$ and the other in $\bar{\mathcal{S}}$, that is,
@@
Conductance (or the Cheeger ratio) of $\mathcal{S}$ is defined to be
@@
A graph can be partitioned into two based on the conductance. This can be accomplished through a sweep over a vector $x\in\mathcal{H}(\mathcal{V})$ for cut (edges), $\partial(\mathcal{S})$, discovery. Suppose $v_{1},\ldots,v_{|\mathcal{V}|}$ is an ordering of vertices such that $x(v_{i})\geq x(v_{i+1})$ for $i=1,\ldots,|\mathcal{V}|-1$. What are known as sweep sets $\mathcal{S}_{j}=\{v_{1},\ldots,v_{j}\}$ for all $j\in\mathcal{N}=\{1,\ldots,|\mathcal{V}|-1\}$ can then be formed. It is the smallest conductance
$\Phi(\mathcal{S}_{*})=\arg\,\min_{j\in\mathcal{N}}\,\Phi(\mathcal{S}_{j})$
which provides the partition.
@@ 2.3 Matrix representation of graphs
Structural information of a graph $\mathcal{G}$ can be encoded in matrix form. Let $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. Connectivity between vertices is captured in the $n\times n$ symmetric adjacency matrix $A$, which has the entries
@@
The degree matrix $D$ is the diagonal matrix with $d_{ii}=\sum_{j=1}^{n}a_{ij}=d(v_{i})$ for $i=1,2,\ldots,n$ and $v_{i}\in V$. A (random walk) transition probability matrix can then be defined as $P=D^{-1}A$.
Edge-vertex connectivity features in the $m\times n$ incidence matrix $B$. For the sole purpose of the definition, an orientation on the graph is required. This entails the specification of an arbitrary but fixed order to the vertices of every edge in $\mathcal{E}$. An edge with ordered vertices is written $\bar{e}=[u,v]$, in which $u=o(\bar{e})$ is the origin vertex and $v=t(\bar{e})$ the terminus vertex. Entries of the incidence matrix are then
$b_{ij}=\begin{cases}1,&\mbox{ if $v_{j}$ is the terminus vertex of $\bar{e}_{i}$, i.e., $v_{j}=t(\bar{e}_{i})$},\\ -1,&\mbox{ if $v_{j}$ is the origin vertex of $\bar{e}_{i}$, i.e., $v_{j}=o(\bar{e}_{i})$},\\ 0,&\mbox{ otherwise}.\end{cases}$
@@
For a connected graph with $n$ vertices, $\text{rank}\,(B)=n-1$.
Proof 2.3.
The following are a result of there being only two non-zero entries, namely $-1$ and $1$, in every row of $B$.
@@
    As each row of $B$ sums to zero, the sum of the columns of $B$ is equal to 00. Therefore, the columns of $B$ are linearly dependent, and $\text{rank}\,(B)<n$. As such, $\text{rank}\,(B)\leq n-1$.
@@
    Consider $Bx=0$. The entries of $x$ must all be the same. Therefore, $\text{rank}\,(B)$ is at least $n-1$.
This leads to the conclusion that $\text{rank}\,(B)=n-1$.
An $m\times m$ matrix $C$ with edge weights on the diagonal, i.e., $C=\text{diag}\left(w(e_{1}),w(e_{2}),\ldots,w(e_{m})\right)$, is often a necessary accompaniment to the incidence matrix.
@@
Note that the Laplacian matrix is symmetric in nature.
Remark 2.4.
Even though the incidence matrix figures in the computation, the Laplacian matrix is independent of orientation.
@@
For a connected graph with $n$ vertices, $\text{rank}\,(L)=n-1$.
Proof 2.6.
Observe that
$\text{rank}\,(B)=\text{rank}\,(C^{-1}CB)\leq\text{rank}\,(CB)\leq\text{rank}\,(B).$
It is thus sufficient to consider the rank of $B^{\top}B$, which is equal to the rank of $B$.
@@ 2.4 Moore-Penrose inverse
Moore in 1920 [9] (at the fourteenth western meeting of the American Mathematical Society) presented an extension to the notion of a nonsingular square matrix inverse that covers (finite-dimensional) rectangular matrices. Even though this extension featured in [23], not much attention seems to have been paid to it. This could possibly be due to the rather unfortunate notation usage [3, 6]. It was not until 1955 that Penrose [30] was to put forward an equivalent theory [6]. Let $Y$ be an $m\times n$ matrix. Following Penrose’s approach, what has become known as the Moore-Penrose inverse is the unique $n\times m$ matrix $Y^{\dagger}$ that satisfies
(1) $YY^{\dagger}Y$ $=Y,$
(2) $Y^{\dagger}YY^{\dagger}$ $=Y^{\dagger},$
(3) $\left(YY^{\dagger}\right)^{*}$ $=YY^{\dagger},$
(4) $\left(Y^{\dagger}Y\right)^{*}$ $=Y^{\dagger}Y,$
where $Y^{*}$ represents the conjugate transpose of $Y$. When $Y$ is real, so is $Y^{\dagger}$, due to the uniqueness of the solution. If $Y$ is nonsingular, then $Y^{\dagger}$ reduces to the familiar inverse of a square matrix, i.e., $Y^{-1}$.
An explicit formula for the Moore-Penrose inverse of $Y$ with full (column) rank can be obtained from [29, Corollary 1 on p. 674] by setting $Y=YI$, in which $I$ is the standard identity matrix:
@@
or $Y^{\dagger}=\left(Y^{\top}Y\right)^{-1}Y^{\top}$, if $Y$ is real.
@@ 2.5 Notation
  • The Hadamard power of an $m\times n$ real matrix $Y$ with respect to a positive integer $p\in\mathbb{Z}_{>0}$ is of the form $Y^{\circ p}=\left(y_{ij}^{p}\right)$. An extension involving $p\in\mathbb{R}$ requires $Y$ having only positive entries.
@@ 3 PageRank problem
Consider $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$. Let $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. Assume a probability vector $r\in\mathcal{H}(\mathcal{V})$, known as the teleportation vector, in which there will be only a single non-zero entry for a starting vertex $s\in\mathcal{V}$, i.e., $r(s)=1$ and $r(v)=0$ for all $v\in\mathcal{V}\setminus\{s\}$. Given a damping factor $\alpha\in(0,1)$, the PageRank vector $x\in\mathcal{H}(\mathcal{V})$ is the solution of the eigenvector problem
(5) $Ux=x\mbox{ with }q^{\top}x=1,$
where $U=\alpha P^{\top}+(1-\alpha)rq^{\top}$. Since $U$ is a stochastic matrix, it follows that the largest eigenvalue is equal to $1$. The PageRank vector, thus, corresponds to the dominant eigenvalue. Moreover, the PageRank vector is a probability vector. This is supported by the normalisation $q^{\top}x=1$.
@@
Introduction of the transition probability matrix definition along with $A=D-L$ yields
$\left(\beta D+L\right)D^{-1}x=\beta r,$
@@
Let $T=\beta D+D^{-1}L$ with $\beta\in(0,1)$. $T^{\text{-1}}$ exists, and $T^{\text{-1}}\geqslant 0$.
Proof 3.2.
Even though the normalised Laplacian matrix is singular in nature (see Proposition 2.5), addition with the scaled degree matrix produces the strictly diagonally dominant matrix $T$, that is,
$|t_{ii}|>\sum_{j=1,j\neq i}^{n}|t_{ij}|$
for $i=1,2,\ldots,n$. Thus, $T$ is nonsingular [13, Corollary 5.6.17], and $T^{\text{-1}}$ exists.
Since $T$ is strictly diagonally dominant with entries $t_{ij}\leq 0$ whenever $i\neq j$ and $t_{ii}>0$ for $1\leq i,j\leq n$, $T$ can be classified as an $M$-matrix according to [32, Property 1.20 on p. 30]. A notable characteristic (among many, refer to [31]) of a nonsingular $M$-matrix is that the entries of the inverse are all non-negative. Subsequently, $T^{\text{-1}}\geqslant 0$.
Attention is now focused on the the system of linear equations
(6) $Tx=\beta r.$
A perspective of $T:\mathcal{H}(\mathcal{V})\to\mathcal{H}(\mathcal{V})$ as an operator can be taken which satisfies
$\left(Tx\right)(v)=\left(\beta d(v)+1\right)x(v)-\frac{1}{d(v)}\sum_{u\sim v}w(u,v)x(u)$
for $v\in\mathcal{V}$. Define
$\tilde{x}(v)=\begin{cases}\left(\beta d(v)+1\right)x(v),&\mbox{if $v\in\mathcal{V}\setminus\{s\}$},\\ \left(\beta d(v)+1\right)x(v)-\beta,&\mbox{if $v=s$}.\end{cases}$
@@
(7) $\tilde{x}(v)=\frac{1}{d(v)}\sum_{u\sim v}w(u,v)x(u).$
That is to say, $\tilde{x}(v)$ for every $v\in\mathcal{V}$ is simply the weighted average of the values at the neighbouring vertices.
@@
Let $\beta\in(0,1)$. Given $\left(Tx\right)(v)=\beta r(v)$ for all $v\in\mathcal{V}$, a maximum value of $x$ can only occur at $s\in\mathcal{V}$ and $x(s)<\frac{1}{d(s)}$.
Proof 3.4.
There are two important implications of (7). A constant $x$ is not feasible, as $x(v)=\tilde{x}(v)\big{/}\left(\beta d(v)+1\right)$ for all $v\in\mathcal{V}\setminus\{s\}$ and
$x(s)=\frac{\tilde{x}(s)}{\beta d(s)+1}+\frac{\beta}{\beta d(s)+1}$
for $s\in\mathcal{V}$. Secondly, a maximum value at any $v\in\mathcal{V}\setminus\{s\}$ cannot be realised. Note that $x(v)\geq 0$ for all $v\in\mathcal{V}$, which is due to $T^{\text{-1}}\geqslant 0$ (see Proposition 3.1) and the non-negative entries of $r$. The maximum value of $x$ can, therefore, only be located at the starting vertex $s\in\mathcal{V}$. Say $x(u)=x(s)$ for all $u\sim s$. Then
$\left(\beta d(s)+1\right)x(s)-\beta<\frac{1}{d(s)}\sum_{u\sim s}w(u,s)x(s)=x(s).$
Clearly, $x(s)<\frac{1}{d(s)}$.

### v2

@@ Abstract
A nonlinear generalisation of the PageRank problem involving the Moore-Penrose inverse of an incidence matrix is developed for local graph partitioning purposes. The Levenberg-Marquardt method with a full rank Jacobian variant provides a strategy for obtaining a numerical solution to the generalised problem. Sets of vertices are formed according to the ranking supplied by the solution, and a conductance criterion decides upon the set that best represents the cluster around a starting vertex. Experiments on both synthetic and real-world inspired graphs demonstrate the capability of the approach to not only produce low conductance sets, but to also recover local clusters with an accuracy that consistently surpasses state-of-the-art algorithms.
@@
[…]
[1]organization=Department of Mathematical Sciences, Aalborg University, addressline=Thomas Manns Vej 23, city=Aalborg Øst, postcode=9220, country=Denmark
[…]
[2]organization=Institute of Computing, Faculty of Informatics, Università della Svizzera italiana, addressline=Via la Santa 1, city=Lugano, postcode=6900, country=Switzerland
@@ 1 Introduction
Graphs are ubiquitous as a means of representing objects and their relationships. It is the very abstraction inherent in the definition of a graph that makes them so widely applicable. Vertices can be any kind of object — such as images [1] or geographical locations [2] — and there are no restrictions on the number of connections between them, which are expressed as edges.
@@
A graph may well exhibit a granular structure. Vertices that share a common property can be uncovered by a process called clustering. The resulting distinct groupings of vertices are referred to as communities or clusters. For example, if all connections are equally valuable, then it may be that communities are distinguished by the relatively high number of intra-community edges compared to the inter-community number of edges.
@@
Spectral clustering is a prominent algorithm that casts clustering as a graph partitioning problem. Clusters are often determined through an analysis of the combinatorial Laplacian matrix eigenvectors [3, 4]. A generalisation of spectral clustering utilising the $p$-Laplacian [5, 6] was pursued in [7]. Improved clustering assignments were demonstrated in the bi-partitioning case. Recursive bisection is, however, required to obtain a higher number of clusters. Direct multiway approaches include approximating the $p$-orthogonality constraint [8, 9], the effective $p$-resistance [10], and exploiting concepts from total variation used commonly in image processing [11].
@@
When only a single cluster around a vertex or vertices is of interest, then an alternative strategy can be employed. Local clustering, as the name suggests, seeks to leverage local structure and information to identify a single cluster. This has the potential to improve algorithmic run-time by avoiding computations over the entirety of a graph.
@@
Treatment of clustering as a partitioning problem carries over into local clustering. A cut can be found using a variation of the PageRank problem [12]. Since a PageRank vector provides a ranking of vertices (refer to [13] for information on functional rankings), sets of vertices can then be assembled according to their order and interrogated to reveal a cut. Similarly to spectral clustering, algorithms based on the $p$-norm have been introduced for local clustering purposes. There is the nonlinear $p$-norm cut algorithm [14], and an algorithm based on the idea of diffusion with $p$-norm network flow [15].
@@
This work takes the system of linear equations that forms the PageRank problem discussed in Section 3 as the starting point in the construction of a local clustering algorithm. Only clusters around a single vertex are considered. A nonlinear generalisation of the PageRank problem defined on a simple, connected and weighted graph that is inspired by the $p$-norm is proposed in Section 4. The Moore-Penrose inverse of the incidence matrix plays a very important role in the generalisation. It is shown in Section 4 that the generalised problem reduces to that of the (linear) PageRank problem in the limit as the number of vertices tends to infinity. Additionally, in Section 4, an infinitesimal perturbation argument offers an insight into the positive effect of the generalised problem on the cluster criterion. The Levenberg-Marquardt method with a full rank Jacobian variant for obtaining a numerical solution to the generalised problem is outlined in Section 5.
@@
The next section presents the theory required for the development of the generalised problem and the identification of a cluster. Section 6 highlights the capability of the proposed algorithm on a number of synthetic and real-world inspired graphs. The algorithm performs strongly on a consistent basis, achieving results better than state-of-the-art algorithms.
@@ 2.1 Notions of graph theory
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set $\mathcal{V}$ of vertices and a set of unordered non-repeating pairs of distinct elements of $\mathcal{V}$ called edges, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Any two vertices $u,v\in\mathcal{V}$ are said to be adjacent or neighbours if they form an edge $e=\{u,v\}$ of $\mathcal{E}$. Their relationship is indicated by $u\sim v$. By definition, there can be no loops, i.e., $\{v,v\}\notin\mathcal{E}$, and only one edge can ever join two vertices. Such a graph is often described as being simple.
If a path exists between every pair of vertices $u$ and $v$ in $\mathcal{G}$, that is, a sequence of vertices $u=u_{1},u_{2},\ldots,u_{n},u_{n+1}=v$ with $e_{i}=\{u_{i},u_{i+1}\}\in\mathcal{E}$ for $i=1,2,\ldots,n$, then the graph is said to be connected.
@@ 2.2 Conductance and sweep-cut
For a subset $\mathcal{S}$ of the vertices in a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$, i.e., $\mathcal{S}\subseteq\mathcal{V}$, the volume of $\mathcal{S}$ has the form
@@
The complement of $\mathcal{S}$ is $\bar{\mathcal{S}}=\mathcal{V}\setminus\mathcal{S}$. When $\mathcal{S}\subset\mathcal{V}$, an edge boundary $\partial(\mathcal{S})$ of $\mathcal{S}$ collects the edges with one vertex in $\mathcal{S}$ and the other in $\bar{\mathcal{S}}$, that is,
@@
Conductance (or the Cheeger ratio) of $\mathcal{S}$ is defined as
@@
A graph can be partitioned into two based on the conductance. This can be accomplished through a sweep over a vector $x\in\mathcal{H}(\mathcal{V})$ for cut (edges), $\partial(\mathcal{S})$, discovery. Suppose $v_{1},\ldots,v_{|\mathcal{V}|}$ is an ordering of vertices such that $x(v_{i})\geq x(v_{i+1})$ for $i=1,\ldots,|\mathcal{V}|-1$. What are known as sweep sets $\mathcal{S}_{j}=\{v_{1},\ldots,v_{j}\}$ for all $j=1,\ldots,|\mathcal{V}|-1$ can then be formed. Let $\mathcal{N}=\{S_{1},\ldots,S_{|\mathcal{V}|-1}\}$. It is the smallest conductance
$\mathcal{S}_{*}=\arg\,\min_{\mathcal{S}\in\mathcal{N}}\,\Phi(\mathcal{S})$
that provides the partition.
@@ 2.3 Matrix representation of graphs
Structural information of a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ can be encoded in matrix form. Let $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. Connectivity between vertices is captured in the $n\times n$ symmetric adjacency matrix $A$, which has the entries
@@
The degree matrix $D$ is the diagonal matrix with $d_{ii}=\sum_{j=1}^{n}a_{ij}=d(v_{i})$ for $v_{i}\in V$ and $i=1,2,\ldots,n$. A (random walk) transition probability matrix can then be defined as $P=D^{-1}A$.
Edge-vertex connectivity features in the $m\times n$ incidence matrix $B$. For the sole purpose of the definition, an orientation on the graph is assumed. This entails the specification of an arbitrary but fixed order to the vertices of every edge in $\mathcal{E}$. An edge with ordered vertices is written $\vec{e}=[u,v]$, in which $u=o(\vec{e})$ is the origin vertex and $v=t(\vec{e})$ the terminus vertex. Entries of the incidence matrix are then
[…]
@@
For a simple and connected graph with $n$ vertices, $\text{rank}\,(B)=n-1$.
Proof.
Based on [16, Lemma 2.2 on p. 12]. The following are a result of there being only two non-zero entries, namely 1-1 and $1$, in every row of $B$.
@@
    As each row of $B$ sums to zero, the columns of $B$ are linearly dependent and $\text{rank}\,(B)<n$. Thus, $\text{rank}\,(B)\leq n-1$.
@@
    Consider $Bx=0$. The entries of $x$ must all be equal due to the connectedness of the graph. As such, $\text{rank}\,(B)$ is at least n1n-1.
In conclusion, $\text{rank}\,(B)=n-1$. ∎
An $m\times m$ matrix $C$ with edge weights on the diagonal, i.e., $C=\text{diag}\left(w(e_{1}),w(e_{2}),\ldots,w(e_{m})\right)$, is often a useful accompaniment to the incidence matrix.
@@
Note that $L$ is symmetric, positive semidefinite and singular [17, Proposition 3.4 on p. 1193].
@@
For a simple, connected and weighted graph with $n$ vertices, $\text{rank}\,(L)=n-1$.
Proof.
It is shown that $\text{ker}\,(L)=\text{ker}\,(B)$ and the rank is provided by Proposition 2.1. If $x\in\text{ker}\,(L)$, then $Lx=0$. Consider $x^{\top}Lx=(Bx)^{\top}C(Bx)=0$. Since $C$ is positive definite, it follows that $Bx=0$ and $x\in\text{ker}\,(B)$. Conversely, if $x\in\text{ker}\,(B)$, then $Bx=0$. Thus, $Lx=0$ and $x\in\text{ker}\,(L)$. ∎
@@ 2.4 Moore-Penrose inverse
Moore in 1920 [18] (at the fourteenth western meeting of the American Mathematical Society) presented an extension to the notion of a nonsingular square matrix inverse that covers (finite-dimensional) rectangular matrices. Even though this extension featured in [19], not much attention seems to have been paid to it. This could possibly be due to the rather unfortunate notation usage [20, 21]. It was not until 1955 that Penrose [22] was to put forward an equivalent theory [21]. Let $Y$ be an $m\times n$ matrix. In Penrose’s approach, $Y^{\dagger}$ is the unique $n\times m$ matrix that satisfies
$YY^{\dagger}Y$ $=Y,$ (1a)
$Y^{\dagger}YY^{\dagger}$ $=Y^{\dagger},$ (1b)
$\left(YY^{\dagger}\right)^{*}$ $=YY^{\dagger},$ (1c)
$\left(Y^{\dagger}Y\right)^{*}$ $=Y^{\dagger}Y,$ (1d)
where $Y^{*}$ represents the conjugate transpose of $Y$. Note that $\text{ker}\,(Y^{\dagger})=\text{ker}\,(Y^{*})$ [21, Theorem 1.2.2 on p. 12]. When $Y$ is real, so is $Y^{\dagger}$, due to the uniqueness of the solution. If $Y$ is nonsingular, then $Y^{\dagger}$ reduces to the familiar inverse of a square matrix, i.e., $Y^{-1}$.
An explicit formula for the Moore-Penrose inverse of $Y$ with full (column) rank can be obtained from [23, Corollary 1 on p. 674] by setting $Y=YI$, in which $I$ is the standard identity matrix:
@@
or $Y^{\dagger}=\left(Y^{\top}Y\right)^{-1}Y^{\top}$, if $Y$ is real. A result that will prove to be useful is recorded in the following proposition.
Proposition 2.3.
For an $m\times n$ matrix $Y$, $(Y^{\dagger}Y)^{\dagger}=Y^{\dagger}Y$.
Proof.
Let $Z=Y^{\dagger}Y$. It follows that $ZZ=(Y^{\dagger}YY^{\dagger})Y=Y^{\dagger}Y=Z$ and $Z^{*}=(Y^{\dagger}Y)^{*}=Y^{\dagger}Y=Z$ by (1b) and (1d), respectively. Say $X=Z$. It is now shown that $X$ satisfies (1a)–(1d):
$ZXZ=ZZZ=ZZ$ $=Z,$
$XZX=ZZZ=Z$ $=X,$
$\left(ZX\right)^{*}=\left(ZZ\right)^{*}=Z^{*}=Z=ZZ$ $=ZX,$
$\left(XZ\right)^{*}=\left(ZZ\right)^{*}=Z^{*}=Z=ZZ$ $=XZ.$
Thus, $X=Z^{\dagger}$. ∎
@@ 2.5 Notation
  • The Hadamard power of an $m\times n$ real matrix $Y$ with respect to a positive integer $p\in\mathbb{Z}_{>0}$ is of the form $^{\circ p}=\left(y_{ij}^{p}\right). YY$ with only positive entries allows for $p\in\mathbb{R}$.
@@ 3 PageRank problem
Consider $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$. Let $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. Assume a probability vector $r\in\mathcal{H}(\mathcal{V})$, known as the teleportation vector, in which there will only be a single non-zero entry for a starting vertex $s\in\mathcal{V}$, i.e., $r(s)=1$ and $r(v)=0$ for all $v\in\mathcal{V}\setminus\{s\}$. Given a damping factor $\alpha\in(0,1)$, the PageRank vector $x\in\mathcal{H}(\mathcal{V})$ is the solution of the eigenvector problem
$Ux=x\mbox{ with }q^{\top}x=1,$ (2)
where $U=\alpha P^{\top}+(1-\alpha)rq^{\top}$. Since $U$ is a stochastic matrix, it follows that the largest eigenvalue is equal to $1$. The PageRank vector, thus, corresponds to this eigenvalue and is a probability vector, as $q^{\top}x=1$.
@@
Introduction of the transition probability matrix definition along with A=DLA=D-L returns
$\left(\beta I+LD^{-1}\right)x=\beta r,$ (3)
@@
Let $T=\beta I+LD^{-1}$ with $\beta\in(0,\infty)$. $T^{-1}$ exists, and $T^{-1}\geqslant 0$.
Proof.
Since all the entries on the main diagonal of $T$ are positive and $TD=\beta D+L=(\beta+1)D-A$ is strictly diagonally dominant, i.e.,
$(\beta+1)d_{ii}>\sum_{j=1,j\neq i}^{n}|a_{ij}|$
for $i=1,2,\ldots,n$, it follows that $T$ is a nonsingular $M$-matrix according to [24, Condition N39 on p. 182]. A notable characteristic of a nonsingular $M$-matrix is that the entries of the inverse are all non-negative. Subsequently, $T^{-1}\geqslant 0$. ∎
A perspective of $T:\mathcal{H}(\mathcal{V})\to\mathcal{H}(\mathcal{V})$ as an operator can be taken, which satisfies
$\left(Tx\right)(v)=\left(\beta+1\right)x(v)-\sum_{u\sim v}\frac{w(u,v)}{d(u)}x(u)$
for each $v\in\mathcal{V}$. Define
$\hat{x}(v)=\begin{cases}\left(\beta+1\right)x(v),&\mbox{if $v\in\mathcal{V}\setminus\{s\}$},\\ \left(\beta+1\right)x(v)-\beta,&\mbox{if $v=s$}.\end{cases}$
@@
$\hat{x}(v)=\sum_{u\sim v}\frac{w(u,v)}{d(u)}x(u).$
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$. Assume $\beta\in(0,\infty)$. Given $\left(Tx\right)(v)=\beta r(v)$, it follows that $x(v)<1$ for all $v\in\mathcal{V}$.
Proof.
A summation of $x(v)$ over all $v\in\mathcal{V}$ yields
$x[\mathcal{V}]$ $=\frac{1}{\beta+1}\sum_{v\in\mathcal{V}}\sum_{u\sim v}\frac{w(u,v)}{d(u)}x(u)+\frac{\beta}{\beta+1}$
$=\frac{1}{\beta+1}\sum_{u\in\mathcal{V}}\frac{x(u)}{d(u)}\sum_{v\sim u}w(u,v)+\frac{\beta}{\beta+1}$
$=\frac{1}{\beta+1}x[\mathcal{V}]+\frac{\beta}{\beta+1}.$
Thus, $x[\mathcal{V}]=1$. Note that $x(v)\geq 0$ for all $v\in\mathcal{V}$, as $T^{\text{-1}}\geqslant 0$ (see Proposition 3.1) and $r(v)\geq 0$ for all $v\in\mathcal{V}$. If $x(v)=1$ and $\hat{x}(v)=\beta+1$ for any $v\in\mathcal{V}\setminus\{s\}$, then by implication $\beta+1=0$. This contradicts $\beta>0$. Similarly, $x(s)=1$ does not hold. As such, $x(v)<1$ for all $v\in\mathcal{V}$. ∎